how many gallons of milk containing 19% butterfat must be mixed with 15 gallons of milk containing 12% butterfat to produce a mixture containing 14% butterfat?
Answer by stanbon(75887)

how many gallons of milk containing 19% butterfat must be mixed with 15 gallons of milk containing 12% butterfat to produce a mixture containing 14% butterfat?
-----------
Equation:
fat + fat = fat
0.19x + 0.12*15 = 0.14(x+15)
--------
19x + 12*15 = 14x + 14*15
5x = 2*15
x = 6 gallons (amt. of 19% milk needed)
---------------------
Cheers,
Stan H.
==================